Please show ALL of your work!

Question 15 Let X denote the number of patients who suffer an infection within a floor of a hospital per month with the following probabilities (See Fig. below) Determine the following probabilities: Round your answer(s) to three decimal places (e.g. 98.765) (a) Less than one infection (b) More than three infections (c) At least one infection (d) No infections 2 PlX-x) 10.69 0.19 0.051 0.069
Answer #1

a)P(less then one infection)=P(X=0)=0.69

b)P(more then 3)=P(X>3)=0

c)P(at least one infection)=1-P(X=0)=1-0.69=0.31

d) P(no infection)=P(X=0)=0.69